OpenStudy (anonymous):

Solve: x^2 − 2x − 24 = 0.

OpenStudy (anonymous):

(x-6)(x+4)=0 x=6 or x=-4

OpenStudy (gorv):

(x-6)*(x+4)=0 x=6,-4

OpenStudy (hexagon001):

(x-6)(x+4)=0 x=6 and x=-4

OpenStudy (anonymous):

Yes, you could use the ones that they gave you but their are more like (x-6)(x+4)=0 x^2=2x+24 or even (x-1)^2-25=0

OpenStudy (anonymous):

x² - 2x - 24 = 0 x² - x = 24 + (- 1)² x² - x = 24 + 1 (x - 1)² = 25 x - 1 = ± 5 = x - 1 - 5, = x - 6 = x - 1 + 5, = x + 4 Answer: (x - 6)(x + 4)

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

your welcome @allie_bear22